# HyperSHAP: Shapley Values and Interactions for Hyperparameter Importance

### v1

@@ Abstract
Hyperparameter optimization (HPO) is a crucial step in achieving strong predictive performance. However, the impact of individual hyperparameters on model generalization is highly context-dependent, prohibiting a one-size-fits-all solution and requiring opaque automated machine learning (AutoML) systems to find optimal configurations. The black-box nature of most AutoML systems undermines user trust and discourages adoption. To address this, we propose a game-theoretic explainability framework for HPO that is based on Shapley values and interactions. Our approach provides an additive decomposition of a performance measure across hyperparameters, enabling local and global explanations of hyperparameter importance and interactions. The framework, named HyperSHAP, offers insights into ablations, the tunability of learning algorithms, and optimizer behavior across different hyperparameter spaces. We evaluate HyperSHAP on various HPO benchmarks by analyzing the interaction structure of the HPO problem. Our results show that while higher-order interactions exist, most performance improvements can be explained by focusing on lower-order representations.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Game-theoretic explanations as defined with HyperSHAP’s hyperparameter importance games can be used to gain insights into hyperparameter values, hyperparameter configuration spaces, datasets, and different hyperparameter optimizers. HyperSHAP can be used for data-specific explanations or explanations across datasets.
@@
Hyperparameter optimization (HPO) is an important step in the design process of machine learning (ML) applications to achieve optimal performance for a given dataset and performance measure (Snoek et al., 2014; Bischl et al., 2023). This is particularly true for deep learning, where hyperparameters describe the neural architecture and steer the learning behavior, e.g., via the learning rate (Zimmer et al., 2021). Also, in the age of generative AI and fine-tuning of large foundation models, HPO is key for achieving state-of-the-art results (Yin et al., 2021; Tribes et al., 2023; Wang et al., 2023).
@@
Hyperparameters affect the generalization performance of models in varied ways, with some having a more significant impact on tuning than others (Bergstra & Bengio, 2012; Hutter et al., 2014; Zimmer et al., 2021). The influence of hyperparameters on generalization performance is highly context-dependent, varying with the dataset characteristics (e.g., size, noise level) and the specific performance measure being optimized (e.g., accuracy, F1 score) (Bergstra & Bengio, 2012; van Rijn & Hutter, 2018). This complexity makes HPO particularly challenging, requiring opaque automated machine learning (AutoML) systems to find optimal configurations within large search spaces (Feurer et al., 2015; Wever et al., 2021). Yet, even after arriving at an optimized hyperparameter configuration, understanding why it outperforms other configurations remains difficult due to intricate effects and interactions among hyperparameters.
Despite their promise, AutoML systems have not fully permeated user groups such as domain experts, ML practitioners, and ML researchers (Lee et al., 2019; Bouthillier & Varoquaux, 2020; Hasebrook et al., 2023; Simon et al., 2023). This lack of adoption stems, in part, from the rigidity of many AutoML systems, which are often difficult to adapt to specific use cases, but is also attributed to the lack of transparency and interpretability (Wang et al., 2019; Drozdal et al., 2020). Studies highlight that a concrete requirement often requested by AutoML system users is interpretability (Wang et al., 2019; Xin et al., 2021; Hasebrook et al., 2023; Sun et al., 2023), and its lack has even led users to favor manual development for high-stakes projects (Xin et al., 2021). For ML researchers, explanations of HPO processes are particularly relevant, as they often prioritize understanding the effectiveness of individual ML components and require control over key aspects of model behavior. Similarly, AutoML researchers need such kind of information to analyze AutoML systems’ performance and behavior. Prior works on hyperparameter importance analysis (Hutter et al., 2014; Watanabe et al., 2023; Theodorakopoulos et al., 2024) and hyperparameter effects (Moosbauer et al., 2021; Segel et al., 2023) show that addressing these interpretability gaps is critical for building trust and enabling more effective use of AutoML systems in a synergetic way with ML experts and data scientists (Lindauer et al., 2024).
1.1 Contribution
In this paper, we formalize HyperSHAP, a post-hoc explanation framework for hyperparameter importance:
@@ 1.1 Contribution
    We define a comprehensive set of 5 explanation games and interpret them using the Shapley value and interactions on three levels: specific configurations, hyperparameter spaces, and optimizer bias.
@@
    With HyperSHAP, we elicit hyperparameter importance and interaction structures for various benchmarks, observing the existence of higher-order interactions.
@@
    We apply HyperSHAP to various explanation tasks and demonstrate its versatility.
1.2 Related Work
Hyperparameter importance (HPI) has gained significant attention in machine learning due to its crucial role in justifying the need for HPO and in attributing performance improvements to specific hyperparameters (Probst et al., 2019; Pushak & Hoos, 2020, 2022; Schneider et al., 2022). A variety of approaches have been developed to assess how different hyperparameters affect the performance of resulting models, ranging from simple (surrogate-based) ablations (Fawcett & Hoos, 2016; Biedenkapp et al., 2017) to sensitivity analyses and eliciting interactions between hyperparameters based on the functional ANOVA framework (Hutter et al., 2014; van Rijn & Hutter, 2018; Bahmani et al., 2021; Watanabe et al., 2023). In this work, we propose a novel approach to quantifying HPI using Shapley values, with a particular focus on capturing interactions between hyperparameters through Shapley interaction indices. We focus on quantifying interactions since, in (Zimmer et al., 2021; Pushak & Hoos, 2022; Novello et al., 2023), it has been noticed that interaction is occasionally comparably low, which could serve as a foundation for a new generation of HPO methods that do not assume interactions to be omnipresent.
Beyond quantifying HPI, to better understand the impact of hyperparameters and the tuning behavior of hyperparameter optimizers, other approaches have been proposed, such as algorithm footprints (Smith-Miles & Tan, 2012), partial dependence plots for hyperparameter effects (Moosbauer et al., 2021) or deriving symbolic explanations (Segel et al., 2023), providing an interpretable model for estimating the performance of a learner from its hyperparameters. In this work, we focus on quantifying the impact of tuning a hyperparameter on the performance.
@@ 2 Hyperparameter Optimization
Hyperparameter optimization (HPO) is concerned with the problem of finding the most suitable hyperparameter configuration (HPC) of a learner for a given task, typically consisting of some labeled dataset $D$ and some performance measure $u$ quantifying the usefulness (Bischl et al., 2023). To put it more formally, let $\mathcal{X}$ be an instance space and $\mathcal{Y}$ a label space and suppose $x\in\mathcal{X}$ are (non-deterministically) associated with labels $y\in\mathcal{Y}$ via a joint probability distribution $P(\cdot\,,\,\cdot)$. Then, a dataset $D=\{(x^{(k)},y^{(k)})\}_{k=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$ is a sample from that probability distribution. Furthermore, a predictive performance measure $u:\mathcal{Y}\times P(\mathcal{Y})\rightarrow\mathbb{R}$ is a function mapping tuples consisting of a label and a probability distribution over the label space to the reals. Given an HPC $\bm{\lambda}\in\bm{\Lambda}$, a learner parameterized with $\bm{\lambda}$ maps datasets $D$ from the dataset space $\mathbb{D}$ to a corresponding hypothesis $h_{\bm{\lambda},D}\in\mathcal{H}:=\{h\mid h:\mathcal{X}\rightarrow P(\mathcal{Y})\}$.
@@
As an HPC $\bm{\lambda}\in\bm{\Lambda}$ typically affects the hypothesis space $\mathcal{H}$ and the learning behavior, it needs to be tuned to the given dataset and performance measure. The task of HPO is then to find an HPC yielding a hypothesis that generalizes well beyond the data used for training. For a dataset $D\in\mathbb{D}$, the following optimization problem needs to be solved:
$\bm{\lambda}^{\ast}\in\underset{\bm{\lambda}\in\bm{\Lambda}}{\arg\max}\int_{(x,y)\sim P(\cdot,\cdot)}u\big{(}y,h_{\bm{\lambda},D}(x)\big{)}\,\,.$
As the true generalization performance is intractable, it is estimated by splitting the given dataset $D$ into training $D_{T}$ and validation data $D_{V}$. Accordingly, we obtain
[…]
$\mathbb{E}_{(D_{T},D_{V})\sim D}\left[\frac{1}{|D_{V}|}\sum_{(x,y)\in D_{V}}u\big{(}y,h_{\bm{\lambda},D_{T}}(x)\big{)}\right].$
@@
Naively, hyperparameter optimization can be approached by discretizing the domains of hyperparameters and conducting a grid search or by a random search (Bergstra & Bengio, 2012). State-of-the-art methods leverage Bayesian optimization and multi-fidelity optimization for higher efficiency and effectiveness (Bischl et al., 2023).
3 Explainable AI (XAI) and Game Theory
@@
Within the field of eXplainable AI (XAI), cooperative game theory has been widely applied to assign contributions to entities, such as features or data points for a given task (Rozemberczki et al., 2022). Most prominently, to interpret predictions of black box models using feature attributions (Lundberg & Lee, 2017) and the Shapley Value (SV) (Shapley, 1953). Shapley Interactions (SIs) (Grabisch & Roubens, 1999) extend the SV by additionally assigning contributions to groups of entities, which reveal synergies and redundancies. Such feature interactions uncover additive structures in predictions, which are necessary to understand complex decisions (Sundararajan et al., 2020). Overall, explanations are summarized by two components (Fumagalli et al., 2024a): First, an explanation game $\nu:2^{\mathcal{N}}\to\mathbb{R}$ is defined as a real-valued set function over the powerset $2^{\mathcal{N}}$ of the $n$ features of interest indexed by $\mathcal{N}=\{1,\dots,n\}$. This explanation game restricts the model prediction to subsets of features and evaluates a property of interest, e.g. the prediction or performance. Second, given an explanation game, interpretable main and interaction effects are constructed using the SV and SIs. Analogously, in Section 4, we define explanation games based on ablations of hyperparameters in $\textsc{Val}_{u}$ and quantify importances with the SV and SIs.
@@
Baseline imputation heavily depends on the choice of baseline (Sturmfels et al., 2020), and marginal and conditional imputation extends the approach to an average over randomized baselines (Sundararajan & Najmi, 2020) as
$\nu_{\mathbf{x}}^{(p)}(S):=\mathbb{E}_{\mathbf{b}\sim p(\mathbf{b})}[f(\mathbf{x}\oplus_{S}\mathbf{b})],$
where $p(\mathbf{b})$ is either the marginal feature distribution or conditioned on $\mathbf{b}=\mathbf{x}\oplus_{S}\mathbf{b}$, respectively. Imputed model predictions define local explanation games that are used to explain predictions of single instances $\mathbf{x}$. In contrast, global explanation games capture derived properties, such as the variance or performance of the imputed model prediction. It was shown that explanations derived from baseline, marginal and conditional imputation are increasingly influenced by the feature distribution $p$ (Fumagalli et al., 2024a).
@@
An explanation game is additively decomposed by the Möbius Interactions (MIs) $m:2^{\mathcal{N}}\to\mathbb{R}$ (Muschalik et al., 2024a), i.e. the Möbius transform (Rota, 1964), for $T\subseteq\mathcal{N}$ as
@@
The MIs describe the pure main and interaction effects, but contain $2^{n}$ non-trivial components in ML applications (Muschalik et al., 2024a). As a remedy, the SV and SIs summarize the MIs into interpretable main and interaction effects of lower complexity. The SV assigns contributions to individuals and is uniquely characterized by four intuitive axioms: linearity (contributions are linear for linear combination of games), symmetry (players with equal contributions obtain equal payout), dummy (players that do not change the payout receive zero payout), and efficiency (the sum of all payouts equals the joint payout). The SV summarizes the MIs as $\phi^{\text{SV}}(i)=\sum_{S\subseteq\mathcal{N}:i\in S}\frac{1}{|S|}m(S)$ for all $i\in\mathcal{N}$. In other words, each MI is equally distributed among the involved players. Yet, the SV does not yield any insights into interactions. Given an explanation order $k\in\{1,\dots,n\}$, the SIs $\Phi_{k}$ extend the SV to assign contributions to subsets of players up to size $k$. For $k=1$ the SIs yield the SV, whereas for $k=n$ the MIs. While there exist multiple variants of SIs, a positive interaction indicates a synergistic effect, whereas a negative interaction indicates redundancies (on average) of the involved features. For instance, the Faithful Shapley Interaction Index (FSII) (Tsai et al., 2023) is defined as the best $k$-additive approximation $\hat{\nu}_{k}(S):=\sum_{L\subseteq S:|L|\leq k}{\Phi_{k}(L)}$ of $\nu(S)$ across all subsets $S$ weighted by the Shapley kernel, cf. Section D.2, which is useful to analyze the degree of interaction. The SIs adjust explanation expressivity and complexity based on practitioner needs, a framework we now parallel in HPO.
4 HyperSHAP: Attributing Importance to Hyperparameters
In hyperparameter optimization (HPO), explanations are needed on different levels of the HPO process, ranging from returned configurations to a qualitative comparison of entire HPO tools. Here, we limit ourselves to four areas, dubbed Ablation, Sensitivity, Tunability, and Optimizer Bias. First, we introduce Ablation (Section 4.1), which we use as the fundamental backbone of HyperSHAP. Based on Ablation, we discuss Sensitivity as an extension of the functional ANOVA framework by Hutter et al. (2014), and compare it theoretically to our novel approach for Tunability (Section 4.2). Tunability is then used to discover Optimizer Bias (Section 4.3), and we conclude with practical aspects of HyperSHAP (Section 4.4). In the following, we let $\mathcal{N}$ be the set of hyperparameters and quantify main and interaction effects based on the SV and SIs of the HPI games. All proofs are deferred to Appendix A.

### v2

# HyperSHAP: Shapley Values and Interactions for Explaining Hyperparameter Optimization

@@ Abstract
Hyperparameter optimization (HPO) is a crucial step in achieving strong predictive performance. Yet, the impact of individual hyperparameters on model generalization is highly context-dependent, prohibiting a one-size-fits-all solution and requiring opaque HPO methods to find optimal configurations. However, the black-box nature of most HPO methods undermines user trust and discourages adoption. To address this, we propose a game-theoretic explainability framework for HPO based on Shapley values and interactions. Our approach provides an additive decomposition of a performance measure across hyperparameters, enabling local and global explanations of hyperparameters’ contributions and their interactions. The framework, named HyperSHAP, offers insights into ablation studies, the tunability of learning algorithms, and optimizer behavior across different hyperparameter spaces. We demonstrate HyperSHAP’s capabilities on various HPO benchmarks to analyze the interaction structure of the corresponding HPO problems, demonstrating its broad applicability and actionable insights for improving HPO.
Code & Appx.https://github.com/automl/HyperSHAP
@@ 1 Introduction
Hyperparameter optimization (HPO) is an important step in the design process of machine learning (ML) applications to achieve strong performance for a given dataset and performance measure (snoek-icml14a; DBLP:journals/widm/BischlBLPRCTUBBDL23). Especially, this is true for deep learning, where hyperparameters describe the architecture and steer the learning behavior (zimmer-tpami21a). Also, for generative AI and fine-tuning of foundation models, HPO is key for achieving the best results (DBLP:conf/acl/YinCSJCL20; llmFinetuningHPO; DBLP:conf/automl/00010A23).
@@
Hyperparameters affect the generalization performance of models in varied ways, with some having a more significant impact on tuning than others (Bergstra-jair12a; fANOVA; zimmer-tpami21a). The impact of hyperparameters on performance is highly context-dependent, varying with the dataset characteristics (e.g., size, noise level) and the specific performance measure being optimized (e.g., accuracy, F1) (Bergstra-jair12a; DBLP:conf/kdd/RijnH18).
This complexity makes HPO particularly challenging, requiring opaque HPO methods to find optimal configurations within large search spaces (feurer-nips15a). Yet, even with an optimized configuration, understanding why it outperforms others remains difficult due to intricate effects and interactions among hyperparameters.
Despite their potential, HPO methods remain underused by domain experts, ML practitioners, and ML researchers (Lee2020; Hasebrook2023; Simon2023). This limited adoption is partly due to their rigidity and poor adaptability to special cases, but also to a lack of interpretability (Wang2019a; Drozdal2020). The latter is a key requirement among HPO users (Wang2019a; Xin2021; Hasebrook2023; Sun2023), and its absence has even led to a shift to manual tuning in high-stakes applications (Xin2021). For ML researchers, explanations are crucial to understand the contribution of individual components and retain control over model behavior. HPO researchers rely on such insights to analyze method performance and behavior. Prior work on hyperparameter importance and effects (fANOVA; DBLP:conf/nips/MoosbauerHCLB21; DBLP:conf/automl/SegelGTBL23; DBLP:conf/ijcai/WatanabeBH23; moo-fANOVA) highlights the need to close interpretability gaps to build trust and foster effective collaboration between HPO tools and ML practitioners (LindauerKKMT0HF24). A complementary view is offered by tunability (tunability), measuring performance gains over defaults to guide whether and what to tune. Yet, explanation methods tailored to tunability remain scarce.
Figure 1: Game-theoretic explanations as defined with HyperSHAP analyze hyperparameter values, hyperparameter spaces, and optimizers. HyperSHAP can be used for data-specific explanations or across datasets.
Contribution
We formalize HyperSHAP, a novel post-hoc HPO-explanation framework:
@@
    We propose HyperSHAP as a collection of 5 explanation games and interpret them using Shapley values and interactions for specific configurations, hyperparameter spaces, and optimizers.
@@
    We showcase how HyperSHAP can be employed for tackling various explanation tasks.
@@
    Comparing to fANOVA, we find that HyperSHAP’s explanations are more actionable to select subsets of hyperparameters for tuning.
  4. (4)
    We provide a publicly available reference implementation of HyperSHAP via GitHub.
2 Related Work
Hyperparameter importance (HPI) has gained significant attention in machine learning due to its crucial role in justifying the need for HPO (DBLP:conf/gecco/PushakH20a; DBLP:journals/telo/PushakH22; DBLP:conf/ppsn/SchneiderSPBTK22), whereas tunability quantifies how much certain hyperparameters can be tuned for specific tasks (tunability). A variety of approaches have been developed to assess how different hyperparameters affect the performance of resulting models, ranging from simple (surrogate-based) ablations (Fawcett-jh16a; DBLP:conf/aaai/BiedenkappLEHFH17) to sensitivity analyses and eliciting interactions between hyperparameters based on fANOVA (fANOVA; DBLP:conf/kdd/RijnH18; elShawiTuneOrNotTune; DBLP:conf/ijcai/WatanabeBH23). In this work, we propose a novel approach to quantifying HPI using Shapley values, with a particular focus on capturing interactions between hyperparameters through Shapley interaction indices. We focus on quantifying interactions since prior works (zimmer-tpami21a; DBLP:journals/telo/PushakH22; DBLP:journals/jscic/NovelloPLC23) noticed that interaction is occasionally comparably low, which could serve as a foundation for a new generation of HPO methods that do not assume interactions to be omnipresent.
Beyond quantifying HPI, to better understand the impact of hyperparameters and the tuning behavior of hyperparameter optimizers, other approaches have been proposed, such as algorithm footprints (DBLP:conf/cec/Smith-MilesT12), partial dependence plots for hyperparameter effects (DBLP:conf/nips/MoosbauerHCLB21) or deriving symbolic explanations (DBLP:conf/automl/SegelGTBL23), providing an interpretable model for estimating the performance of a learner from its hyperparameters.
@@ 3 Hyperparameter Optimization
Hyperparameter optimization (HPO) is concerned with the problem of finding the most suitable hyperparameter configuration of a learner for a given task, typically consisting of some labeled dataset $D$ and some performance measure $u$ quantifying the usefulness (DBLP:journals/widm/BischlBLPRCTUBBDL23). To put it formally, let $\mathcal{X}$ be an instance space and $\mathcal{Y}$ a label space and suppose $x\in\mathcal{X}$ are (non-deterministically) associated with labels $y\in\mathcal{Y}$ via a joint probability distribution $\mathbb{P}$.
Then, a dataset $D=\{(x^{(k)},y^{(k)})\}_{k=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$ is a sample from that probability distribution. Furthermore, a predictive performance measure $u:\mathcal{Y}\times P(\mathcal{Y})\rightarrow\mathbb{R}$ is a function mapping tuples consisting of a label and a probability distribution over the label space to the reals. Given a configuration $\bm{\lambda}\in\bm{\Lambda}$, a learner parameterized with $\bm{\lambda}$ maps datasets $D$ from the dataset space $\mathbb{D}$ to a corresponding hypothesis $h_{\bm{\lambda},D}\in\mathcal{H}:=\{h\mid h:\mathcal{X}\rightarrow P(\mathcal{Y})\}$.
@@
As a configuration $\bm{\lambda}\in\bm{\Lambda}$ typically affects the hypothesis space $\mathcal{H}$ and the learning behavior, it needs to be tuned to the given dataset and performance measure. The task of HPO is then to find a configuration yielding a hypothesis that generalizes well beyond the data used for training. For a dataset $D\in\mathbb{D}$, the following optimization problem needs to be solved: $\bm{\lambda}^{\ast}\in\underset{\bm{\lambda}\in\bm{\Lambda}}{\arg\max}\int_{(x,y)\sim\mathbb{P}}u\big(y,h_{\bm{\lambda},D}(x)\big).$ As the true generalization performance is intractable, it is estimated by splitting the given dataset $D$ into training $D_{T}$ and validation data $D_{V}$. Accordingly, we obtain $\bm{\lambda}^{\ast}\in\underset{\bm{\lambda}\in\bm{\Lambda}}{\arg\max}\,\textsc{Val}_{u}(\bm{\lambda},D)$, with
$\textsc{Val}_{u}(\bm{\lambda},D):=\mathbb{E}_{(D_{T},D_{V})\sim D}\sum_{(x,y)\in D_{V}}\frac{u\big(y,h_{\bm{\lambda},D_{T}}(x)\big)}{|D_{V}|}.$
@@
Naïvely, HPO can be approached by discretizing the domains of hyperparameters and conducting a grid search or by a random search (Bergstra-jair12a). More commonly, state-of-the-art methods often leverage Bayesian optimization and multi-fidelity optimization for higher efficiency and effectiveness (DBLP:journals/widm/BischlBLPRCTUBBDL23).
4 Explainable AI and Game Theory
@@
Within the field of eXplainable AI (XAI), cooperative game theory has been widely applied to assign contributions to entities, such as features or data points for a given task (DBLP:conf/ijcai/RozemberczkiWBY22). Most prominently, it is used to interpret predictions of black-box models using feature attributions (Lundberg.2017) and the Shapley Value (SV) (Shapley.1953). Shapley Interactions (SIs) (Grabisch.1999) extend the SV by additionally assigning contributions to groups of entities, which reveal synergies and redundancies.
Such feature interactions reveal additive structures essential for understanding complex predictions (Sundararajan.2020). Explanations consist of two components (Fumagalli.2024a): (1) an explanation game $\nu:2^{\mathcal{N}}\to\mathbb{R}$, a set function over feature subsets of the $n$ features of interest indexed by $\mathcal{N}=\{1,\dots,n\}$ that evaluates properties such as prediction or performance; (2) interpretable main and interaction effects derived from the SV and SIs. Analogously, the next section defines explanation games over hyperparameter ablations in $\textsc{Val}_{u}$ , using the SV and SIs to quantify tunability.
@@
Baseline imputation is highly sensitive to the chosen baseline (sturmfels2020visualizing). Marginal and conditional imputation extend this by averaging over randomized baselines (Sundararajan.2020b): $\nu_{\mathbf{x}}^{(p)}(S):=\mathbb{E}_{\mathbf{b}\sim p(\mathbf{b})}[f(\mathbf{x}\oplus_{S}\mathbf{b})],$ where $p(\mathbf{b})$ is the marginal or conditional feature distribution. The imputed predictions define local games for explaining individual predictions, while global games capture aggregate properties, e.g., variance or performance. As such, explanations increasingly reflect the underlying distribution $p$ (Fumagalli.2024a).
@@
An explanation game is additively decomposed by the Möbius Interactions (MIs) $m:2^{\mathcal{N}}\to\mathbb{R}$ (Muschalik.2024), i.e. the Möbius transform (rota1964foundations), for $T\subseteq\mathcal{N}$ as
@@
The MIs capture pure main and interaction effects but contain $2^{n}$ non-trivial components, too many for practical interpretation in ML applications (Muschalik.2024). To reduce this complexity, the SV and SIs summarize the MIs into interpretable effects. The SV assigns contributions to individuals, is uniquely characterized, and satisfies four axioms: linearity, symmetry, dummy, and efficiency. The SV summarizes the MIs distributing each MI among the involved players via $\phi^{\text{SV}}(i)=\sum_{S\subseteq\mathcal{N}:i\in S}\frac{1}{|S|}m(S)$ for all $i\in\mathcal{N}$. Yet, the SV does not uncover interactions. Given an explanation order $k\in\{1,\dots,n\}$, the SIs $\Phi_{k}$ extend the SV to assign contributions to subsets of players up to size $k$. For $k=1$ the SIs yield the SV and the MIs for $k=n$. Various forms of SIs exist, where positive values indicate synergy and negative values signal redundancy among the involved features. For instance, the Faithful Shapley Interaction Index (FSII) (Tsai.2022) defines the best $k$-additive approximation $\hat{\nu}_{k}(S):=\sum_{L\subseteq S:|L|\leq k}{\Phi_{k}(L)}$ weighted by the Shapley kernel, enabling quantification of interaction strength. SIs thus offer a flexible trade-off between expressivity and complexity, a framework we now adapt to HPO.
5 Explaining Hyperparameter Optimization
Explanations in HPO are needed at multiple levels, from individual configurations to qualitative comparisons of HPO tools. Here, we consider four areas, dubbed Ablation, Sensitivity, Tunability, and Optimizer Bias. We begin with Ablation as the foundation of HyperSHAP, extend it to Sensitivity (showing links to fANOVA by fANOVA), and compare it theoretically to Tunability. Tunability then serves to uncover Optimizer Bias. We conclude with practical considerations of HyperSHAP. Let $\mathcal{N}$ denote the set of hyperparameters; we quantify main and interaction effects based on the SV and SIs of the explanation games. Proofs are deferred to the appendix.